# Why do explanations fail? A typology and discussion on failures in XAI

### v1

@@ Introduction
This causes the explanations to be potentially perceived by users as being incomplete or incorrect (e.g., lack of knowledge or lack of transparency on the XAI method). As discussed in the previous section, users may thus draw on external information and cognitive processes to interpret the explanations, leading to potentially one or more XAI failures. For instance, users may expect that there can be intuitive changes in counterfactual examples because they have experienced the same logical path in real life (Suffian et al. 2022).
@@
We believe that XUIs allow to better organize, complete and display ML explanations according to the user needs (Chromik and Butz 2021). Moreover, accounting for the users is key to design such interfaces, which thus forces AI practitioners to adopt a user-centered approach when conceiving ML systems (see Section Observation: Some failures result from the interaction between components of the explanation process.). Previous studies have demonstrated the usefulness of visual interfaces to present ML explanations (e.g., (Szymanski, Millecamp, and Verbert 2021; Ooge, Kato, and Verbert 2022)) but there are other interaction modalities, in particular the conversational mode. As users have progressively become more familiar with conversational AIs (e.g. ChatGPT), the provided explanations may thus be presented in conversational interfaces and become a real dialog, i.e. questions from the users and corresponding answers from the machine (Miller 2019). This could open new perspectives for design principles such as introducing a narrative logic that allows temporizing the current information. For instance (Nourani et al. 2021) suggest to control primary information to mitigate negativity bias. We believe that studying such a modality for the display of explanations would allow to better understand users’ processes for analyzing and understanding ML explanations.
Conclusion
In this work, we have proposed a typology of XAI failures allowing to understand why failures happen, how to mitigate them, and how to distinguish them from one another. We believe it can help AI developers and designers better understand XAI systems and their limitations. Leveraging this typology, we have identified promising research avenues for XAI. Besides these directions, future works include investigating the potential combinations of multiple failures and their consequences on user understanding. Furthermore, a better understanding of each failure could be leveraged to formally propose a diagnostic framework to help identifying their origins in the explanation process.

### v2

@@ 6.3 Towards More XUI
This causes the explanations to be potentially perceived by users as incomplete or incorrect (e.g., lack of knowledge or lack of transparency on the XAI method). As discussed in Sect. 5.3, users may thus draw on external information and cognitive processes to interpret the explanations, potentially leading to XAI failures. For instance, users may expect that there can be intuitive changes in counterfactual examples because they have experienced the same logical path in real life (Suffian et al. 2022).
@@
We believe that XUIs allow to better organize, complete and display ML explanations according to the user needs (Chromik and Butz 2021). Moreover, accounting for the users is key to design such interfaces, which thus forces AI practitioners to adopt a user-centered approach when conceiving ML systems (see Sect. 6.1). Previous studies have demonstrated the usefulness of visual interfaces to present ML explanations (see e.g., Szymanski, Millecamp, and Verbert (2021); Ooge, Kato, and Verbert (2022)) but there are other interaction modalities, in particular the conversational mode. As users have progressively become more familiar with conversational AIs (e.g. ChatGPT), the provided explanations may be presented in conversational interfaces and become a dialog, i.e. questions from the users and corresponding answers from the machine (Miller 2019; Ribera and Lapedriza 2019). This could open new perspectives for design principles such as introducing a narrative logic that allows temporizing the current information. For instance primary information can be controlled to mitigate the negativity bias (Nourani et al. 2021). We believe that studying such a modality for the display of explanations would allow to better understand users’ processes for analyzing and understanding ML explanations.
7 Conclusion
In this work, we have proposed a typology of XAI failures allowing to understand why failures happen, how to mitigate them, and how to distinguish them from one another. We believe it can help AI developers and designers better understand XAI systems and their limitations. Leveraging this typology, we have identified promising research avenues for XAI. In addition to these directions, future works will include investigating the potential interaction between multiple failures. Besides better understanding their connections to one another, the co-occurrence or superposition of several failures raises crucial questions regarding their consequences on user understanding. Furthermore, a better understanding of each failure and these interactions could be leveraged to formally propose a diagnostic framework to help identifying their origins in the explanation process.